OpenStudy (anonymous):

The quadrilateral has vertexes at (-1, 7), (-6, 1), (4, 1) and (-1, -5). Its sides are all the same length. How long is each side?

jimthompson5910 (jim_thompson5910):

Use the distance formula to find the distance from (-1,7) to (-6,1)

OpenStudy (anonymous):

Thank you :), the distance formula is d=sqrt(x2-x1)^2+(y2-y1)^2 right?

jimthompson5910 (jim_thompson5910):

The distance formula is d = sqrt( (x2 - x1)^2 + (y2 - y1)^2 ) In your case, (x1,y1) = (-1,7) (x2,y2) = (-6,1) So you plug all that in to get d = sqrt( (x2 - x1)^2 + (y2 - y1)^2 ) d = sqrt( (-6 - (-1))^2 + (1 - 7)^2 ) d = sqrt( (-6 +1)^2 + (1 - 7)^2 ) and I'll let you finish up. Tell me what you get.

jimthompson5910 (jim_thompson5910):

You are correct

jimthompson5910 (jim_thompson5910):

So what distance do you get?

OpenStudy (anonymous):

sqrt61 ?

jimthompson5910 (jim_thompson5910):

yep each side is \(\Large \sqrt{61}\) units long

OpenStudy (anonymous):

thank you so much :)

jimthompson5910 (jim_thompson5910):

yw